OpenStudy (anonymous):

(2m^3)^0(3m^6)^-1 its simplifying,the answer is 1/3m^6 i just dont know how to get that

OpenStudy (anonymous):

\[(2m^3)^0*(3m^6)^{-1}\] Ok we know that anything to the power of 0 =1 so for instance \(x^0=1\) Hence we can conclude that \((2m^3)^0=1\) Another property is that any number to the power of a negative value means *flip me* so \(x^{-1}=\large\frac{1}{x}\) or \(x^{-2}=\large\frac{1}{x^2}\) so therefore \[(3m^6)^{-1}=\frac{1}{3m^6}\] Therefore we have \[1*\frac{1}{3m^6}=\frac{1}{3m^6}\]

OpenStudy (anonymous):

wait why does it say "math processing error"

OpenStudy (anonymous):

wtf is wrong with os???

OpenStudy (anonymous):

http://puu.sh/7jmmL.png

OpenStudy (anonymous):

Take a look at that link ^

OpenStudy (anonymous):

oh ok thnkss

OpenStudy (anonymous):

ooo i get it now thanxx